k3nn3dyj:

Which expression is equivalent to one over threep + 15? one over three(p + 45) one over three(p + 5) 3(p + 45) 3(p + 5)

AP:

Original equation: \[\frac{ 1 }{ 3 }p+15\] options: A.) \[\frac{1}{3}p+45\] B.) \[\frac{1}{3}(p+5)\] C.) \[3(p+45)\] D.) \[3(p+5)\]

Mercury:

old question but will respond so this can be closed take (1/3)p + 15 and factor out 1/3